# FreshStack: Building Realistic Benchmarks for Evaluating Retrieval on Technical Documents

### v1

@@ 4.3. Retrieval Evaluation Metrics
$\alpha$-nDCG@k. Introduced by Clarke et al. (2008), this variant of Normalized Discounted Cumulative Gain (nDCG) measures search diversification. The $\alpha$ parameter is a geometric penalization for redundant documents, i.e., each redundant document achieves a penalization of ×$\times(1-\alpha)$. Despite the metric being used for different user intents, we utilize it to ensure document rankings reference diverse nuggets in the answer.
@@
5. Pooling & Qualitative Evaluation
Using FreshStack, we construct five IR datasets focused on document retrieval evaluation with a subset of question-answer pairs. [^8] For each question in the FreshStack dataset, the curated answer is present, allowing us to extend answer evaluation in RAG as potential future work. In this section, we attempt to answer RQ2 by evaluating techniques that contribute to judgment pools. Next, we calibrate the quality of GPT-4o against an expert human annotator. We first evaluate the retrieval baselines during nugget-level support judgment (or sampling pools) with inference and oracle settings.
In FreshStack, we are constructing a test evaluation dataset. Therefore, we can use the Stack Overflow answer or its variants in constructing judgment pools, as discussed previously in Section 3.4. We pool and sample documents from different settings and techniques, similar to how existing question answering datasets are constructed, such as Natural Questions (Kwiatkowski et al., 2019) or XOR-TyDI (Asai et al., 2021), which assess relevance by calculating the answer overlap in the document.
Experimental Settings. We perform retrieval with four techniques and baselines (explained in Section 3.4) and an ensemble fusion of the top 100 documents, with each model score normalized and summed up. Evaluation metrics include $\alpha$-nDCG@10, Coverage@20, and Recall@50. We use GPT-4o with a temperature setting of 0.1999Separately, we tested temperatures of 0.1 and 0.7, observing an identical downstream retrieval accuracy during FreshStack construction. for both the automatic stages. Nugget generation uses a grading notes prompt with the question and answer, and support assessment uses chain-of-thought prompting (Wei et al., 2022), judging up to a maximum of 20 documents simultaneously with a list of nuggets generated for each question. Finally, we sample and judge the top 20 fusion documents from each technique and setting, including the question to avoid sampling holes, highlighting the importance of document diversity in our judgment pools.
5.1. Pooling Results
We outline the results achieved on document judgment pools during FreshStack dataset construction with four different sampling techniques and models, including inference and oracle settings. Table 4 shows results for all five topics. Key takeaways and findings are discussed below:
Overall Highlights. Table 4 reveals two key findings: (1) Techniques in the Oracle setting significantly outperform techniques from the inference setting. We observe that both the Stack Overflow answer and nuggets techniques help pool documents relevant to the question, and (2) Fusion outperforms all individual models, highlighting the value of diversity in model choice, aiding in our judgment pools.
Inference Setting. GPT-4o Sub-Questions achieves the best pooling results for four topics (except Yolo v7 & v8), showing that decomposing a Stack Overflow question into smaller sub-questions helps pool documents relevant to the question. GPT-4o closed-book answer excels only on Yolo v7 & v8, which we suspect due to possible data contamination. These overall confirm that FreshStack contains non-contaminated data, as GPT-4o closed-book answer (For example, the ChatGPT generated answer) cannot retrieve relevant documents for answering niche questions in the FreshStack dataset on most topics.
Oracle Setting. Stack Overflow Nuggets achieves the best results on all topics (except Laravel 10 & 11), showing that GPT-4o-generated nuggets help pool more documents relevant to the question, instead of using the Stack Overflow answer. Amongst the individual models, BM25 achieves the best $\alpha$-nDCG@10 on all topics in both oracle methods (except Godot4), which asserts the importance of lexical approaches required during the construction of judgment pools in FreshStack.
5.2. Qualitative Analysis
In our work, a crucial component is the automatic construction of nuggets and nugget-level document judgments with GPT-4o. To assess the LLM quality, we calibrate using an expert annotator (ML researcher) conducting a qualitative analysis on a small subset of LangChain, evaluating the quality of generated nuggets and nugget-document support labels for 60 randomly sampled questions.
Nugget Quality Judgment Quality
Precision 90.1 % Relevant 71.7 %
Recall 96.6 % Partially Relevant 11.7 %
Groundedness 96.4 % Non-Relevant 16.6 %
Table 5. Expert evaluation of GPT-4o nugget quality and nugget-document relevance judgments on LangChain.
5.2.1. Nugget Quality Evaluation
For nugget quality evaluation, we ask the annotator to answer the following questions (A, B, and C) after reading the Stack Overflow question, answer, and list of nuggets. (1) $A$ does the nugget produce hallucinated content? requiring a boolean response (2) $B$ is the information provided in the nugget minor or redundant? also requiring a boolean response. After annotating all nuggets, we ask (3) $C$ how many additional nuggets are required to cover all key ideas, requiring an integer in the response.
Evaluation Metrics. After answering all the above questions, we measure quality by calculating three metrics: (1) precision: whether nuggets generated are accurate, (2) recall: whether nuggets cover the key aspects of the answer and (3) groundedness: whether nuggets produce non-hallucinated content, i.e., outside the scope of the answer. More formally, we define them as follows:
(2) $\text{Precision}=\dfrac{|Nuggets|-\mathrm{sum}(B)}{|Nuggets|}$
(3) $\text{Recall}=\dfrac{|Nuggets|-\mathrm{sum}(B)}{|Nuggets|-\mathrm{sum}(B)+C}$
(4) $\text{Groundedness}=\dfrac{|Nuggets|-\mathrm{sum}(A)}{|Nuggets|},$
where $|Nuggets|$ denotes the count of nuggets for a given question.
Experimental Results. As shown in Table 5, nuggets in the topic of LangChain achieve above 90% in precision and 96% in recall and groundedness, indicating GPT-4o can generate high-quality nuggets required in the FreshStack framework. Most nuggets are well-grounded, i.e., do not produce hallucinated content (3.6% error), and cover the key aspects of the answer in terms of recall (3.4% error). Precision errors are higher (9.9% error), showing nuggets may contain either minor or repeated information. Within these errors, the last positioned nugget is not informative in almost 50% of the time, and either the first or second positioned nugget in the rest of the error cases.
5.2.2. Relevance Judgment Quality Evaluation
We assess the relevance between nuggets and documents in nugget-level support. Since judging all documents (including negatives) for each nugget is cumbersome with a limited budget, we qualitatively evaluate the relevant pairs. We sample one positive document per question, totaling 60 randomly sampled nugget-document pairs. The annotator labels the relevance on a three-level scale: relevant, partially relevant, and non-relevant.
Figure 3. Main experiment results with retrieval and reranker baselines (including fusion) on five topics. Best scores in the Oracle setting are taken from Table 4 (Stack Overflow Nuggets in blue, Stack Overflow Answer in red). The reranker is the Voyage AI rerank-2 model (AI, 2024a). Plots show $\alpha$-nDCG-10, Coverage20, and Recall@50 scores for each topic (top to bottom).
6. Main Experiments
In this section, we evaluate leading neural retrievers and rerankers for the document retrieval setting on the FreshStack dataset, addressing RQ3 posed in our introduction. All models are evaluated out-of-the-box, i.e., using only the Stack Overflow question to retrieve documents, and do not include any information about the Stack Overflow answer or nuggets, ensuring a fair assessment.
Experimental Settings. We evaluate the retrieval models used as baselines during pooling in FreshStack: BM25, BGE (Gemma-2), E5 Mistral 7B, Voyage-large-2, and the ensemble fusion. In addition, we evaluate the Voyage AI rerank-2 (AI, 2024a) as the reranker with a 16K context length, reranking the top 50 documents retrieved from the first-stage retrieval system. Metrics used for evaluation are defined in Section 4.3: $\alpha$-nDCG@10, Coverage@20, and Recall@50.
6.1. Model Evaluation Results
Figure 3 shows the results for each retrieval baseline without and with VoyageAI rerank-2 reranking queries on each topic. Each model category is color-coded. We plot the $\alpha$-nDCG@10, Coverage@20, and Recall@50 scores from left to right. Key takeaways and findings are discussed below:
Accuracy gap between Oracle indicates plenty of headroom. Techniques from the Oracle setting (using Stack Overflow answers or nuggets) achieve a substantially higher $\alpha$-nDCG@10, Coverage@20, and Recall@50 in contrast to all models, including ensemble fusion and reranking with VoyageAI rerank-2. This highlights the complexity of answering FreshStack questions and demonstrates the headroom for improvement in existing code-based retrieval models to decrease the gap. Further research should explore techniques to improve models on specialized topics, containing challenging and long questions in the FreshStack dataset, instead of overfitting on existing academic IR benchmarks such as BEIR (Thakur et al., 2021).
Ensemble fusion outperforms individual models. Individual retrieval models demonstrate limited success on the FreshStack dataset; whereas, the ensemble fusion of four retrieval models outperforms each retrieval model across all metrics ($\alpha$-nDCG@10, Coverage@20, and Recall@50) and all five topics, except $\alpha$-nDCG@10 on Godot4. This highlights a crucial point in retrieval systems: a compound retrieval system (Zaharia et al., 2024), developed as an ensemble of retrieval models or something similar, is required to retrieve documents for niche and challenging topics, as an individual retriever is not sufficient, at present. While the ensemble fusion of retrieval systems achieves the best scores on the FreshStack dataset, their inefficiency at inference time adds up due to individual model inference, requiring alternatives.
Opportunities to improve reranking. When using weak first-stage retrieval, neural rerankers typically improve document ranking (Thakur et al., 2021), although it has been recently shown that this is not always the case when strong first-stage retrieval is used (Zamani et al., 2022; Jacob et al., 2024). Consistent with these recent observations, reranking provides benefits over BM25 for all topics in the FreshStack dataset. However, for our dense retrievers, reranking provides a clear benefit on some but not all datasets. Specifically, while the reranker enhances $\alpha$-nDCG@10, Coverage@20, and Recall@50 for LangChain, Yolo v7 & v8, and Godot4, it reduces those metrics on Laravel 10 & 11 and Angular 16, 17 & 18 for the dense retrievers. We suspect the reranker is better in certain programming languages such as Python, and we keep it as future work to understand the limitations of the neural reranker (Jacob et al., 2024). Separately, we also tried GPT-4o-mini (OpenAI, 2024) naively as a zero-shot listwise reranker following Sun et al. (2023) and found preliminary results underperformed heavily. We suspect this was due to formatting challenges when concatenating multiple long documents, and we leave the exploration of LLMs as rerankers (perhaps with fine-tuning) to future work.
7. Discussion
@@
FreshStack is a general framework for building challenging IR datasets. We apply the framework to community-sourced questions (with curated answers) and documents sourced from GitHub repositories. The framework is adaptable to other domains like Stack Exchange or internal forums. While we benchmark a few key retrieval models, future work would continue to benchmark code-focused retrieval models like voyage-3 (AI, 2024b), Code-T5+ (Wang et al., 2023), CodeRankEmbed (Suresh et al., 2024), and Jina-Code-v2 (Günther et al., 2023), and rerankers such as CodeRankLLM (Suresh et al., 2024).
@@
Answer Evaluation. We focused on the evaluation of the retrieval setting primarily due to two reasons: (1) Existing RAG datasets evaluate retrieval using relevance criteria only, however, we evaluated models based on both diversity and relevance criteria, and (2) a crucial step in FreshStack is sourcing and building a document corpus and developing a general framework for high-quality pools and automatic judgments, which we can evaluate better in the retrieval setting. As we have curated Stack Overflow answers available, future work can extend on evaluating the quality of LLM-based answer generation with metrics such as nugget-based recall, which calculates how many nuggets are supported within a system’s response. Given the high difficulty of retrieval, validated with results observed in Section 6, we anticipate answer generation by LLM to be similarly challenging.
@@
Benchmark Contamination. The FreshStack dataset builds on Stack Overflow data and is susceptible to future data contamination. To address this, FreshStack is intentionally designed to be generic and extensible to new domains and languages. To mitigate data contamination, the framework can add newer questions in existing topics, retire old and contaminated topics, and add newer topics that develop in the future. The relevance of FreshStack in the community relies on a continued commitment to keeping it updated in the upcoming years.
8. Conclusion
Acknowledgements.
We thank Sean Kulinski and Alexander Trott for helping us set up the grading notes prompt required in nugget generation. We also thank Jacob Portes, Max Marion, Matei Zaharia, and others from Databricks who provided feedback at the early stages of the project.

### v2

@@ C.1 Retrieval Evaluation Metrics
$\alpha$-nDCG@k. Introduced by Clarke et al. [9], this variant of Normalized Discounted Cumulative Gain (nDCG) measures search diversification. The $\alpha$ parameter is a geometric penalization for redundant documents, i.e., each redundant document achieves a penalization of ×$\times(1-\alpha)$. Despite the metric being used for different user intents, we utilize it to ensure document rankings reference diverse nuggets in the answer. We would ask the reader to refer to  Clarke et al. [9] for more information.
@@
C.2 RAG Evaluation Metric
All Strict ($A_{strict}$)
Introduced in Pradeep et al. [52], for each query, we have a list of nuggets generated from Section 3.3, and for each RAG answer generated, we have a record of which nuggets it contains, in terms of a three-way judgment: support, partial_support, and no_support. The final step is to compute the score for the RAG answer to a query $q$. The score of a run is simply the mean of the score for each query. We compute the following scores per query:
We calculate a score based on all nuggets in the RAG answer, but with strict nugget matching. For a given nugget $i$:
$p_{i}$ =$=\begin{cases}1&\text{if assignment = {support}}\\ 0&\text{otherwise}\end{cases}$ (3)
The “All Strict” score is then calculated as:
$A_{strict}=\frac{\sum_{i}p_{i}}{|Nuggets|},$
where $|Nuggets|$ denotes the count of nuggets for a given query $q$.
Appendix D FreshStack Instance Description
Each FreshStack dataset instance contains the following four components, as shown in Figure 1. A complete example of a dataset instance is shown in Table 8.
  • Relevance Judgments: Unlike traditional IR benchmarks, such as BEIR [68], which contain question and document-level relevance judgments, FreshStack datasets contain nugget-level relevance judgments for document chunks.
Appendix E Discussion and Future Work
@@
FreshStack is a holistic framework for building challenging IR and RAG evaluation datasets. We apply the framework to community-sourced questions (with curated answers) and documents sourced from GitHub repositories. The framework is adaptable to other domains like Stack Exchange or internal forums.
While we benchmarked a few key retrieval models in our work, we would like to benchmark code-focused retrieval models like voyage-3 [2], Code-T5+ [76], CodeRankEmbed [66], and Jina-Code-v2 [19], and rerankers such as CodeRankLLM [66], in the future.
@@
Answer Evaluation. We focused on the evaluation of the retrieval setting primarily due to two reasons: (1) existing RAG datasets evaluate retrieval using relevance criteria only, however, we evaluated models based on both diversity and relevance criteria, and (2) a crucial step in FreshStack is sourcing and building a document corpus and developing a general framework for high-quality pools and automatic judgments, which we can evaluate better in the retrieval setting. We evaluated the quality of LLM-based answer generation with nugget-based recall with $A_{strict}$ metric [51, 52], which calculates how many nuggets are supported within a system’s response. However, we keep an in-depth evaluation of the RAG answer, accounting for other factors such as fluency or support [67], as future work.
@@
Benchmark Contamination. The FreshStack dataset is built on Stack Overflow data, making it susceptible to future data contamination. Newer released LLMs such as GPT-4.1 with a recent knowledge cutoff date [^9] https://help.openai.com/en/articles/9624314-model-release-notes of June 2024, provide the possibility of dataset contamination with GPT-4.1 as FreshStack queries originated from January 2023 until June 2024 (as shown in Figure 3). To mitigate data contamination, the FreshStack framework can add newer questions in existing topics, retire old and contaminated topics, and add newer topics that develop in the future. The relevance of FreshStack in the community relies on a continued commitment to keeping it updated in the upcoming years.
Figure 3: Timeline versus frequency of how many FreshStack queries were asked on Stack Overflow in every quarter. All queries included in FreshStack were asked between January 2023 and June 2024, with the highest frequencies observed in 2024, showing the growing importance of all five topics.
Figure 4: Token distribution of Stack Overflow questions and answers for all topics in FreshStack. Unlike other benchmarks, FreshStack questions (highlighted in green) are much longer than their answers (highlighted in maroon).
Topic GitHub Repositories & Licenses
LangChain [1] langchain-ai/langchain: MIT license
[2] langchain-ai/langchainjs: MIT license
[3] langchain-ai/langchain-nextjs-template: MIT license
[4] chroma-core/chroma: Apache-2.0 license
[5] openai/openai-cookbook: MIT license
[6] openai/openai-python: Apache-2.0 license
[7] run-llama/llama_index: MIT license
[8] Azure-Samples/openai: MIT license
[9] Azure-Samples/azure-search-openai-demo: MIT license
[10] huggingface/transformers: Apache-2.0 license
Yolo v7 & v8 [1] ultralytics/ultralytics: AGPL-3.0 license
[2] ultralytics/docs: AGPL-3.0 license
[3] pytorch/pytorch: Modified BSD license
[4] WongKinYiu/yolov7: GPT-3.0 license
[5] opencv/opencv: Apache-2.0 license
Laravel 10 & 11 [1] laravel/framework: MIT license
[2] laravel/laravel: MIT license
[3] laravel/laravel.com: MIT license
[4] laravel/docs: MIT license
[5] laravel/breeze: MIT license
[6] livewire/livewire: MIT license
[7] php/php-src: PHP license
[8] php/doc-en: PHP license
[9] php/web-php: PHP license
Angular 16, 17 & 18 [1] angular/angular: MIT license
[2] angular/components: MIT license
[3] angular/angular-cli: MIT license
[4] microsoft/TypeScript: Apache-2.0 license
Godot4 [1] godotengine/godot: MIT license
[2] godotengine/godot-demo-projects: MIT license
[3] godotengine/godot-docs: CC BY 3.0 license
[4] godotengine/godot-website: MIT license
[5] GDQuest/learn-gdscript: MIT license
[6] dotnet/csharplang: GPL license
Table 7: A list of the GitHub repositories with their licenses for constructing the document collection for each topic in FreshStack.
LangChain Query ID: 78256389
Stack Overflow Query Title: Chromadb from_documents function giving error. Text: The following function was working till a few days ago but now gives this error: ValueError: Expected EmbeddingFunction._call_ to have the following signature: odict_keys([’self’, ’input’]), got odict_keys([’args’, ’kwargs’]) Please see https://docs.trychroma.com/embeddings for details of the ‘EmbeddingFunction‘ interface. Please note the recent change to the ‘EmbeddingFunction‘ interface: https://docs.trychroma.com/migration#migration-to-0416---november-7-2023. I am not sure what changes are necessary to work with this. 1def create_chromadb(link): 2 embedding_function = SentenceTransformerEmbeddings(model_name="all-MiniLM-L6-v2") 3 loader = TextLoader(link) 4 documents = loader.load() 5 6 # Split the documents into chunks (no changes needed here) 7 text_splitter = RecursiveCharacterTextSplitter(chunk_size=1000, chunk_overlap=500) 8 chunks = text_splitter.split_documents(documents) 9 10 # Update for new EmbeddingFunction definition 11 # D is set to the type of documents (Text in this case) 12 D = Union[str, List[str]] # Adjust based on your document format (single string or list of strings) 13 embedding_function: EmbeddingFunction[D] = embedding_function 14 15 # Initialize Chroma with the embedding function and persist the database 16 db = Chroma.from_documents(chunks, embedding_function, ids=None, collection_name="langchain", persist_directory="./chroma_db") 17 db.persist() 18 print(f"Saved {len(chunks)} chunks") 19 20 return db
Stack Overflow Answer I slightly modify your code, using ‘HuggingFaceEmbeddings‘ instead of ‘SentenceTransformerEmbeddings‘. 1from langchain_community.embeddings import HuggingFaceEmbeddings 2embedding = HuggingFaceEmbeddings(model_name="sentence-transformers/all-MiniLM-L6-v2") 3 4from langchain_community.vectorstores import Chroma 5db = Chroma.from_documents(documents=chunks, embedding=embedding, persist_directory="/tmp/chroma_db") 6db.persist()
GPT-4o Nuggets 1. The error is due to a mismatch in the function signature expected by ‘Chroma.from_documents‘ when using ‘SentenceTransformerEmbeddings‘. 2. Use ‘HuggingFaceEmbeddings‘ instead of ‘SentenceTransformerEmbeddings‘ to resolve the error. 3. Import ‘HuggingFaceEmbeddings‘ from ‘langchain_community.embeddings‘. 4. Initialize ‘HuggingFaceEmbeddings‘ with the model name “sentence-transformers/all-MiniLM-L6-v2”. 5. Pass the initialized ‘HuggingFaceEmbeddings‘ to the ‘Chroma.from_documents‘ function.
Retrieved GitHub Document Document ID: langchain/templates/intel-rag-xeon/ingest.py_0_1486 Supported Nuggets: Nugget #2, Nugget #4, and Nugget #5. 1import os 2 3from langchain.text_splitter import RecursiveCharacterTextSplitter 4from langchain_chroma import Chroma 5from langchain_community.document_loaders import UnstructuredFileLoader 6from langchain_community.embeddings import HuggingFaceEmbeddings 7from langchain_core.documents import Document 8 9def ingest_documents(): 10 """ 11 Ingest PDF to Redis from the data/ directory that 12 contains Edgar 10k filings data for Nike. 13 """ 14 # Load list of pdfs 15 data_path = "data/" 16 doc = [os.path.join(data_path, file) for file in os.listdir(data_path)][0] 17 18 print("Parsing 10k filing doc for NIKE", doc) 19 text_splitter = RecursiveCharacterTextSplitter(chunk_size=1500, chunk_overlap=100, add_start_index=True) 20 loader = UnstructuredFileLoader(doc, mode="single", strategy="fast") 21 chunks = loader.load_and_split(text_splitter) 22 23 print("Done preprocessing. Created", len(chunks), "chunks of the original pdf") 24 25 # Create vectorstore 26 embedder = HuggingFaceEmbeddings(model_name="sentence-transformers/all-MiniLM-L6-v2") 27 28 documents = [] 29 for chunk in chunks: 30 doc = Document(page_content=chunk.page_content, metadata=chunk.metadata) 31 documents.append(doc) 32 33 # Add to vectorDB 34 _ = Chroma.from_documents(documents=documents, collection_name="xeon-rag", embedding=embedder, persist_directory="/tmp/xeon_rag_db") 35 36 37if __name__ == "__main__": 38 ingest_documents()
Table 8: A complete example of a dataset instance from LangChain in FreshStack. The relevant nuggets supported by the retrieved GitHub document are highlighted in green.